Castles in the Sand generates a rate of return of 10% on its investments and maintains a plowback ratio of 0.30. Its earnings this year will be $4 per share. Investors expect a rate of return of 8% on the stock.

a. Find the price and P/E ratio of the firm. (Do not round intermediate calculations. Round your answers to 2 decimal places.) Price P/E ratio

b. Find the price and P/E ratio of the firm if the plowback ratio is reduced to 0.20. (Do not round intermediate calculations. Round your answers to 2 decimal places.) Price P/E ratio
Answer #1

The growth rate can be calculated as:

= ROE* b

= 0.1 * 0.3

= 3%

So, the price of stock is :

= D1/ Re - g

= $4* 0.7/ 0.08 - 0.03

= $56

So, the P/E ratio is:

= $56/ 4

= 14

b. If the plowback is reduced to 0.2, then the growth rate is :

= 0.1* 0.2

= 2%

So, the price of stock is:

= $4* 0.7/ 0.08 - 0.02

=$46.67

So, P/E IS :

= $46.67 / 4

= 11.67 ( Rounded off to two decimal places)